Elemental analysis of a 3.64 mg sample of unknown hydrocarbon shows that it contains 3.167396 mg of carbon. If the molecular formula is C9Hx, what is the value of x?
Answer #1

The value of x is 18.

Mass of sample=3.64mg

Mass of carbon = 3.167mg

Mass of hydrogen = 3.64-3.167= 0.473mg

Moles of carbon=3.167/12×10^3=2.369×10^-4

Moles of hydrogen=0.473/1×10^3=4.73×10^-4

Ratio is determined by diving with smallest number

For carbon = 2.369×10^-4/2.369×10^-4 = 1

For hydrogen=4.73×10^-4/2.369×10^-4=1.8=2

Empirical formula =CH2

Molecular formula = n×empirical formula

Thus C9Hx=n × CH2

thus n = 9

Hence x=9×2=18

Thus answer is x=18.